# Integer quadratic forms and extensions of subsets of linearly independent roots

### v2

@@ 5.1. $D$-components $\mathscr{L}_{D_{l+1}}(D_{l})$ and $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$
(ii) The linkage label vectors $\varphi^{\nabla}$ and $-\delta^{\nabla}$ coincide.
@@
$(\gamma,\mu_{max}+\tau_{l+1})=\begin{cases}2\text{ for }\gamma=\tau_{l+1},\text{ since }\mu_{max}\perp\tau_{l+1},\\ 0\text{ for }\gamma=\tau_{l},\text{ since }(\tau_{l},\mu_{max})=1,(\tau_{l},\tau_{l+1})=-1,\\ 0\text{ for }\gamma=\tau_{i},\text{ where }i<l.\end{cases}$ (5.5)
@@
(iii) The number of roots of $\varPhi(D_{l})$ is $2l(l-1)$, see [Bo02, Table IV]. Then by (4.7) and (4.8) we have
$\mid\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))\mid~{}=~{}\mid\mathscr{L}_{D_{l+1}}(D_{l})\mid~{}\leq~{}2(l+1)l-2l(l-1)=4l.\$
@@
[…]
@@
$1$ $\begin{array}[]{ccccccc}2&3&4&3&2&1&0\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&3&2\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&0&0&-1\\ &0\end{array}$
$2$ $\begin{array}[]{ccccccc}2&3&4&3&2&1&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&3&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&0&-1&1\\ &0\end{array}$
$3$ $\begin{array}[]{ccccccc}2&3&4&3&2&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&-1&1&0\\ &0\end{array}$
$4$ $\begin{array}[]{ccccccc}2&3&4&3&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&-1&1&0&0\\ &0\end{array}$
$5$ $\begin{array}[]{ccccccc}2&3&4&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&-1&1&0&0&0\\ &0\end{array}$
$6$ $\begin{array}[]{ccccccc}2&3&5&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&5&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}-1&1&0&0&0&0\\ &-1\end{array}$
$7$ $\begin{array}[]{ccccccc}2&3&5&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}2&4&5&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}-1&0&0&0&0&0\\ &1\end{array}$
@@ 5.3. The size and structure of linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$
$\mid\mathscr{L}_{E_{6}}(D_{5}(a_{1}))\mid~{}=~{}\mid\mathscr{L}_{E_{6}}(D_{6})\mid~{}\leq~{}\mid\varPhi(E_{6})\mid-\mid\varPhi(D_{5})\mid~{}=~{}72-40=32.\$
@@
$\mid\mathscr{L}_{E_{7}}(D_{6}(a_{2}))\mid~{}=~{}\mid\mathscr{L}_{E_{7}}(D_{6}(a_{1}))\mid~{}=~{}\mid\mathscr{L}_{E_{7}}(D_{6})\mid~{}\leq~{}\mid\varPhi(E_{7})\mid-\mid\varPhi(D_{6})\mid~{}\leq~{}126-60=66.\$
@@
$\mid\mathscr{L}(D_{6})\mid~{}=~{}\mid\mathscr{L}(D_{6}(a_{2}))\mid~{}=~{}\mid\mathscr{L}(D_{6}(a_{1}))\mid~{}\leq~{}66+12=78.\$
@@
$\begin{array}[]{cccccc}\tau_{6}&\tau_{1}&\tau_{2}&\tau_{3}&\tau_{4}&\tau_{5}\\ &&\beta_{2}\\ \end{array}$
Consider maximal and minimal roots ±$\pm\mu$ in $E_{7}$. Here, ±$\pm\mu\in\varPhi(E_{7})\backslash\varPhi(D_{6})$ and ±$\pm\mu_{max}\not\in\varPhi(D_{6})$. Note that ±$\pm\mu$ are orthogonal to any simple root except for $\tau_{6}$. Then vectors ±$\pm\mu^{\nabla}$ are zero, i.e.,
$\mid\mathscr{L}(D_{6})\mid~{}=~{}\mid\mathscr{L}(D_{6}(a_{1}))\mid~{}=~{}\mid\mathscr{L}(D_{6}(a_{2}))\mid~{}\leq~{}78-2=76.\$ (5.8)
@@
$\mid\mathscr{L}_{E_{8}}(D_{7}(a_{2}))\mid~{}=~{}\mid\mathscr{L}_{E_{8}}(D_{7}(a_{1}))\mid~{}=~{}\mid\mathscr{L}_{E_{8}}(D_{7})\mid~{}\leq~{}\mid\varPhi(E_{8})\mid-\mid\varPhi(D_{7})\mid~{}\leq~{}240-84=156.\$
@@
$\mid\mathscr{L}_{E_{8}}(D_{7}(a_{2}))\mid~{}=~{}\mid\mathscr{L}_{E_{8}}(D_{7}(a_{1}))\mid~{}=~{}\mid\mathscr{L}_{E_{8}}(D_{7})\mid~{}\leq~{}156-14=142.\$
@@
Figure 10. For $l>4$, the partial linkage system $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$ contains $2l$ linkages. For $l>7$, $\mathscr{L}(D_{l}(a_{k}))=\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$
@@ A.1. Gabrielov transformations
Let $B=B_{\Gamma}$ be the (partial) Cartan matrix associated with the diagram $\Gamma$, $E_{ij}$ be the elementary $n\times n$ matrix with unique nonzero entry a $1$ at slot $(i,j)$, and $G^{B}_{ij}$ the linear transformation given by matrix $G^{B}_{ij}=I-b_{ij}E_{ij}$. For each $i\neq j$ in $\{1,\dots,n\}$, the Gabrielov transformation $\mathscr{G}_{ij}$ is defined as the following composition:

### v3

@@ 5.1. $D$-components $\mathscr{L}_{D_{l+1}}(D_{l})$ and $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$
(ii) The linkage label vectors $\varphi^{\nabla}$ and $\delta^{\nabla}$ coincide.
@@
[…]
@@
(iii) The number of roots of $\varPhi(D_{l})$ is 2l(l1)2l(l-1), see [Bo02, Table IV]. Then by (4.7) and (4.8) we have
$\mid\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))\mid~=~\mid\mathscr{L}_{D_{l+1}}(D_{l})\mid~\leq~2(l+1)l-2l(l-1)=4l.\\$
@@
Vectors $\eta_{i}+\lambda_{i}$ from (5.6) are orthogonal to any $\tau_{i},\;2\leq i\leq 8$, so $\eta_{i}^{\nabla}=-\lambda_{i}^{\nabla}$. Note that there is no need orthogonality for $\tau_{1}$, see (1.15). Further, we calculate the label vectors $\eta_{i}^{\nabla}$ (or $\lambda_{i}^{\nabla}$). They are listed in the last column of Table 5.2.
@@
11 $\begin{array}[]{ccccccc}2&3&4&3&2&1&0\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&3&2\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&0&0&-1\\ &0\end{array}$
22 $\begin{array}[]{ccccccc}2&3&4&3&2&1&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&3&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&0&-1&1\\ &0\end{array}$
33 $\begin{array}[]{ccccccc}2&3&4&3&2&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&-1&1&0\\ &0\end{array}$
44 $\begin{array}[]{ccccccc}2&3&4&3&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&-1&1&0&0\\ &0\end{array}$
55 $\begin{array}[]{ccccccc}2&3&4&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&-1&1&0&0&0\\ &0\end{array}$
66 $\begin{array}[]{ccccccc}2&3&5&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&5&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}-1&1&0&0&0&0\\ &-1\end{array}$
77 $\begin{array}[]{ccccccc}2&3&5&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}2&4&5&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}-1&0&0&0&0&0\\ &1\end{array}$
@@ 5.3. The size and structure of linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$
$\mid\mathscr{L}_{E_{6}}(D_{5}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{6}}(D_{6})\mid~\leq~\mid\varPhi(E_{6})\mid-\mid\varPhi(D_{5})\mid~=~72-40=32.\\$
@@
$\mid\mathscr{L}_{E_{7}}(D_{6}(a_{2}))\mid~=~\mid\mathscr{L}_{E_{7}}(D_{6}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{7}}(D_{6})\mid~\leq~\mid\varPhi(E_{7})\mid-\mid\varPhi(D_{6})\mid~\leq~126-60=66.\\$
@@
$\mid\mathscr{L}(D_{6})\mid~=~\mid\mathscr{L}(D_{6}(a_{2}))\mid~=~\mid\mathscr{L}(D_{6}(a_{1}))\mid~\leq~66+12=78.\\$
@@
[…]
Consider maximal and minimal roots $\pm\mu$ in $E_{7}$. Here, $\pm\mu\in\varPhi(E_{7})\backslash\varPhi(D_{6})$ and $\pm\mu_{max}\not\in\varPhi(D_{6})$. Note that $\pm\mu$ are orthogonal to any simple root except for $\tau_{6}$. Then vectors $\pm\mu^{\nabla}$ are zero, i.e.,
$\mid\mathscr{L}(D_{6})\mid~=~\mid\mathscr{L}(D_{6}(a_{1}))\mid~=~\mid\mathscr{L}(D_{6}(a_{2}))\mid~\leq~78-2=76.\\$ (5.8)
@@
$\mid\mathscr{L}_{E_{8}}(D_{7}(a_{2}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7})\mid~\leq~\mid\varPhi(E_{8})\mid-\mid\varPhi(D_{7})\mid~\leq~240-84=156.\\$
@@
$\mid\mathscr{L}_{E_{8}}(D_{7}(a_{2}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7})\mid~\leq~156-14=142.\\$
@@
Figure 10. For $l>4$, the partial linkage system $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$ contains $2l$ linkages. For $l>7$, $\mathscr{L}(D_{l}(a_{k}))=\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$ .
@@ A.1. Gabrielov transformations
Let $B=B_{\Gamma}$ be the (partial) Cartan matrix associated with the diagram $\Gamma$, $E_{ij}$ be the elementary $n\times n$ matrix with unique nonzero entry a $1$ at slot $(i,j)$, and $G^{B}_{ij}$ - the linear transformation given by matrix $G^{B}_{ij}=I-b_{ij}E_{ij}$. For each $i\neq j$ in $\{1,\dots,n\}$, the Gabrielov transformation $\mathscr{G}_{ij}$ is defined as the following composition: